OpenStudy (flowey15):

How long is the arc subtended by an angle of 7π/5 radians on a circle of radius 25 cm A) 7π B) 35π C) 125π D) 175π

OpenStudy (jameshorton):

Just change the form of them or solve them by using the same formula. Find the reference to angle A = - 15 Pi / 4 i.e The given angle is not positive and less than 2Pi. We can use the positive and less than 2Pi coterminal Ac to angle A. Ac = - 15 Pi / 4 + 2 (2 Pi) = Pi / 4 Angle A and Ac are coterminal and have the same reference angle. Ac is in quadrant I, therefore A r = A c = Pi / 4 ANOTHER WAY Degree measure of angle is based upon the in a circle and radian measure is based upon as another way to describe one complete circle. Radian measure of an angle is the length of the arc intercepted on a circle of radius 1 by an angle in standard position on a coordinate plane. Or equivalently, the radian measure of a central angle in standard position on a coordinate plane is the ratio of the intercepted arc length to the radius of the circle We can generalize the correspondence between degree and radian measure since we know that 180º corresponds to radians. Using an equal sign to describe this relationship (the radians and degrees are not actually “equal” measurements) we can write: We can use these ratios to convert radian measure into degrees and vice versa.

OpenStudy (flowey15):

that's all it says

OpenStudy (jameshorton):

did u get it

OpenStudy (faiqraees):

Arc length = Angle * radius Arc length = 7π/5 * 25 Arc Length = 35π